OpenStudy (anonymous):

Calculate the surface area of a cone with radius of 4 in. and slant height of 10 in

OpenStudy (msmr):



OpenStudy (msmr):

actually, you don't need to know h, sorry. I thought you were looking for volume.

OpenStudy (msmr):

to find the SA of the cone, you need to find the area of the base and the area of the other part. area of base is just the area of a circle, so πr^2 = 4^2π = 16π area of the other part =πrs, where s = slant height, so 4*10π

OpenStudy (msmr):

16π + 40π = 56π